# Decoding Gray Matter: large-scale analysis of brain cell morphometry to inform microstructural modeling of diffusion MR signals

### v2

@@ 3.3 Cellular shape reference values
All Cell Types Only Neurons Only Glia
$R_{soma}\ (\mu m)$ 2 29absent229$\geq$ 2\ -\ 29≥ 2 - 29 $\geq 7$ $\geq 8$ $\geq 4.0$
$\langle R_{branch}\rangle_{s}\ (\mu m)$ 0.3 4absent0.34$\geq$ 0.3\ -\ 4≥ 0.3 - 4 $\geq 1.6$ $\geq 1.6$ $\geq 1.6$
$CV_{branch}$ 0.3 5.80.35.80.3\ -\ 5.80.3 - 5.8 $1.2$ $1$ $1$
$L_{branch}\ (\mu m)$ 5 14551455\ -\ 1455 - 145 $54$ $64$ $11$
$S/V_{branch}\ (\mu m^{-1})$ 4 164164\ -\ 164 - 16 $8$ $8$ $8$
$\langle\mu OD_{branch}\rangle_{s}$ 0.1 0.50.10.50.1\ -\ 0.50.1 - 0.5 $0.3$ $0.3$ $0.3$
$\langle R_{c}\rangle_{s}(\mu m)$ 4 475absent4475$\geq$ 4\ -\ 475≥ 4 - 475 $\geq 55$ $\geq 64$ $\geq 18$
$\tau_{branch}$ 0.80 0.950.800.950.80\ -\ 0.950.80 - 0.95 $0.88$ $0.90$ $0.92$
$BO$ 4 194194\ -\ 194 - 19 $9$ $9$ $13$
$N_{proj}$ 2 112112\ -\ 112 - 11 $6$ $5$ $9$
$\eta_{soma}\ (\%)$ 1 86absent186$\leq$ 1\ -\ 86≤ 1 - 86 $\leq 18$ $\leq 14$ $\leq 31$
$R_{domain}\ (\mu m)$ 40 7564075640\ -\ 75640 - 756 $334$ $400$ $48$
Table 3: Reference values for all the morphological features of neuronal and glial cells. The ranges and mean values obtained from the whole dataset investigated are reported, together with mean values for only neurons and glia. The $\geq$’ and $\leq$’ are used when the estimated value of the corresponding feature may be slightly (on average $<20\%$) under- or over-estimated, respectively, given the known limitations of the approach used.
Shape
Cell Type Species Watson parameter Orientation dispersion Fractional Anisotropy
Mouse/Rat 0.99 0.50 $0.21\pm 0.11$
Microglia Monkey 1.7 0.34 $0.21\pm 0.11$
Human n.a. n.a. n.a.
Mouse/Rat 2.1 0.28 $0.34\pm 0.21$
Astrocyte Monkey n.a. n.a. n.a.
Human n.a. n.a. n.a.
Mouse/Rat 2.1 0.28 $0.24\pm 0.17$
Oligodendrocyte Monkey n.a. n.a. n.a.
Human n.a. n.a. n.a.
Mouse/Rat 1.9 0.31 $0.33\pm 0.19$
Pyramidal Monkey 2.0 0.30 $039\pm 0.25$
Human 0.67 0.62 $0.45\pm 0.18$
Mouse/Rat 5.8 0.11 $0.75\pm 0.16$
Granule Monkey n.a. n.a. n.a.
Human 3.4 0.18 $0.70\pm 0.15$
Mouse/Rat 5.3 0.12 $0.74\pm 0.08$
Purkinje Monkey n.a. n.a. n.a.
Human n.a. n.a. n.a.
Mouse/Rat 2.1 0.28 $0.25\pm 0.19$
Basket Monkey 0.89 0.54 $0.39\pm 0.28$
Human 0.98 0.51 $0.21\pm 0.21$
Mouse/Rat 1.1 0.47 $0.31\pm 0.16$
Gabaergic Monkey n.a. n.a. n.a.
Human 2.4 0.25 $0.25\pm 0.14$
Mouse/Rat 1.0 0.50 $0.24\pm 0.13$
Glutamatergic Monkey n.a. n.a. n.a.
Human 2.7 0.23 $0.74\pm 0.04$
@@ 3.4 Distribution of some structural features of interest for dMRI modelling
The full distributions of values for some structural features of interest, obtained by merging together all the estimated values from all the species for each cell type, are shown in Fig.6. Given the increasing interest of the microstructure imaging through dMRI community in estimating glia microstructure and the striking difference between the neuronal components of cerebral and cerebellar cortices (e.g., Purkinje cells only in cerebellum), we decided to group them into three classes: glia, cortical neurons and cerebellar neurons, and use three different colors to simplify the visualization of the results.
Figure 6: Probability density functions of soma radius, branch length, intra-branch residence time and intra-cellular residence time for each cell types and all the species. The residence times were calculated from the surface to volume ratios with a permeability of 20 $\mu m/s$ [57] and diffusivity typical of water.
@@ 3.5 Topological distance between cell types
The persistence maps of all the cell types for all species together are shown in Fig.7. These persistence maps are used for each species and each cell type to estimate cell-type specific topological distances, shown in Fig.8.
In rodent cells, there is a notably high global topological distance between glial cells and neurons. This indicates that glial cells exhibit a fundamentally different topological organization compared to neurons. This difference in topology becomes even more evident when considering the comparison along the same length scale. Since glial cells are significantly smaller in size, they inherently display a much smaller topological persistence. This size-related constraint underscores their distinct spatial and structural properties relative to neurons.
Within rodent neurons, granule cells stand out by exhibiting a higher topological distance compared to other neuronal types. This disparity is likely related to their characteristically low branch order, which reflects their simpler dendritic structures and reduced connectivity compared to more complex neurons.
In human cells, however, no significant or consistent trends in topological distance are observed, suggesting greater variability or less pronounced differences in cellular topology among the various cell types.
When examining hominid cells, which include both monkey and human samples, microglia demonstrate a higher topological distance compared to neurons. This observation reinforces the trend seen in rodent cells, further supporting the notion that glial cells, maintain a distinct and conserved topological profile across different species.
In the cross-species comparison between rodent and hominid cells, microglia consistently show a high topological distance relative to neurons, mirroring the trend observed within each species. Additionally, the comparitively low topological distance for microglia across species suggest that their topological properties are consistent, pointing to a shared structural and functional organization in microglia across species.
@@ 4.1 A Quantitative View of Gray Matter Microstructure
GM Intra-Cellular Space. There is currently a lack of in-depth morphological analysis of brain-cell structures of relevance for modelling water diffusion in the GM intra-cellular space. In this work we propose a first analysis with the aim to fill this gap. We estimated a comprehensive set of morphological features useful to GM microstructure modelling from reconstructions of microscopy data from three species. We estimated that neural soma size ranges from  2 to  30 $\mu$m in radius with an average of  7 $\mu$m and surface-to-volume ratio S/V 0.5 $\mu$m-1. Neurons have on average soma twice as big as glial cells, fewer projections radiating from the soma and less projection coverage of soma surface. The radius of cellular projections ranges from  0.3 to  4 $\mu$m with average value 1.6 $\mu$m and S/V 8 $\mu$m-1, similar between neurons and glia. Cellular projections’ microscopic orientation dispersion, as defined in [43], is  0.10-0.50, with average value  0.30, similar between neurons and glia; curvature radius is  4-475 $\mu$m, with average value of 55 $\mu$m. On average, neuronal projections have a curvature radius  6 times larger than glial projections. The branching order of neural cell is  4-20, with average value of 9. Glial cells have branching order  1.5 times larger than neurons. The tortuosity of the branch of neural cell projections is  1.05-1.27, with average value of  1.14, similar between neurons and glial cells. The projections of neural cells extend to distances of 40-756 $\mu$m, with neurons on average 400 $\mu$m and glial cells on average 50 $\mu$m. For completeness, we also report on the relevant features of the GM extra-cellular space and neural cell membrane permeability to water from the literature.
Neural Cell Membrane Permeability to Water. Several works suggest water exchange between unmyelinated neurites and ECS and/or soma occurring on time scales comparable to typical dMRI clinical acquisitions, i.e.  10-100 ms. Although there is not a consensus yet, some works report exchange times $t_{ex}$ for ex vivo mouse brain  5-10 ms [58, 59, 30], suggesting membrane permeability  125 $\mu$m/s (like red blood cells), others report $t_{ex}$ for in vivo mouse brain of  20-40 ms [57, 29, 60], suggesting membrane permeability  2-35 $\mu$m/s. Using the latter estimates of membrane permeability, and our estimates of the cellular and branch surface-to-volume ratios we estimated distribution of intra-domain and intra-branch residence times that span from a few milliseconds to tens of milliseconds (Fig.6). On average, glial cells have shorter residence times than neurons, and small neurons like granule cells have much shorter residence times than large neurons like pyramidal cells. These observations support the idea of a wide distribution of exchange times within the GM, with water within glial cells in faster exchange with ECS than that within neurons (10-20 ms versus 20-50 ms). We have to acknowledge here that in our analysis we did not include dendritic spines, boutons and glial leflets, that can occupy up to $20\%$ of the GM volume [61]. Recent studies have demonstrated that diffusion-mediated exchange between cellular projections’ shaft and these small lateral protrusions is another mechanism of water exchange, different from the permeative one, but happening on similar time scales (from a few millisecond to a few tens of milliseconds) [62, 63, 64]. Therefore, the interplay between diffusion-mediated and permeative exchange cannot be neglected when analysing and intepreting time-dependent dMRI measurements in the GM. Appropriate modelling and acquisition strategies to disentangle the two mechanisms and provide unbiased estimates are still missing, and represent an exciting avenue for future research in the field.
@@ 4.2 General Implications for Biophysical Modelling
Here we provide a few illustrative examples demonstrating how this study can inform biophysical modelling. Assuming an intra-cellular diffusivity D=2 $\mu m^{2}/ms$ representative of water and 0.40 $\mu m^{2}/ms$ representative of intracellular brain metabolites, and considering a typical single diffusion encoding acquisition with gradient pulse duration/separation delta/delta and diffusion time $td<60\ ms$, we can infer from Tab.3 that:
@@
  • Soma restriction can be significant for both water and metabolites: given the low $\eta_{soma}$, the soma can significantly restrict diffusion when (from [65]) $5Dt_{d}\geq R_{soma}^{2}$ , that is for $t_{d}\geq 5\ ms$ for water and $t_{d}\geq 25\ ms$ for metabolites, given a soma radius $R_{soma}\approx 7\ \mu m$. This is supported by previous findings which demonstrate the impact of soma contribution on the final dMRI signal [28, 41]
@@
  • Cellular domain restriction is generally negligible for both water and metabolites: given the size of the cellular domain, it will only significantly restrict molecular diffusion when $5Dt_{d}\geq R_{domain}^{2}$, that is for $t_{d}\geq 160\ ms$ for water and $t_{d}\geq 800\ ms$ for metabolites, given $R_{domain}\geq 40\ \mu m$, which is far longer than typically used diffusion times $t_{d}$.
@@
  • Impact of projections curvedness is generally negligible for both water and metabolites: the impact of curvedness can only be significant when (from [66] $2D\Delta,2D\delta\geq(<R_{c}>_{s})^{2}$, that is for $\Delta,\delta\geq 750\ ms$ for water and $\Delta,\delta\geq 3750\ ms$ for metabolites, given $<R_{c}>_{s}\approx 55\ \mu m$, which is far longer than the values typically. This is further supported by [41].
@@
  • Impact of projection undulation can be significant for both water and metabolites: the values of $<\mu OD_{branch}>_{s}$ estimated from the real cellular data match those simulated in [43], from which it can be concluded that undulation can bias the estimation of projection radius in grey matter. However, the average branch radius $<R_{branch}>_{s}\approx 1.6\ \mu m$ is far below the resolution limit of conventional water dMRI techniques [67].
@@
  • Impact of branching is generally negligible for both water and metabolites: given the branch length $L_{branch}\approx 54\mu m$, the exchange between branches is negligible for $t_{d}<<L_{barnch}^{2}/(2D)\approx 750\ ms$ for water and $t_{d}<<3750\ ms$ for metabolites, significantly longer than typical $t_{d}$ used. Further supported by [39, 41].
@@
  • Impact of membrane permeability for cellular projections can be significant for water. The average intra-branch residence times given the estimated $S/V_{branch}$ are $\approx 20\ ms$ for neurons and $\approx 10\ ms$ for glial cells. The permeative exchange is thus negligible only for $t_{d}<20\ ms$ for neuronal projections and for $t_{d}<10\ ms$ for glial projections; which is not satisfied in conventional water dMRI applications. Further supported by [29, 41, 60].
@@
  • Impact of projections orientation dispersion can be significant. Cells with highly oriented and polarized projections, such as Purkinje cells, granule cells and pyramidal neurons have high FA ($>0.50$) and low orientation dispersion ($<0.25$); while the projections of most of the glial cells and other neuronal cells have low FA ($<0.50$) and high orientation dispersion ($>0.25$). This supports growing evidence that FA, e.g., from DTI, and orientation dispersion estimates, e.g. from NODDI, can discriminate different cytoarchitectural domains (or layers) in hippocampus, cortex and cerebellum [68, 69, 70, 71, 72].

### v4

@@ 3.4 Cellular shape reference values
All Cell Types Only Glia Only Neuron
$R_{domain}$ $\geq 8\ -\ 750$ $240$ $60$ $300$
$N_{proj}$ $\geq 1\ -\ 13$ 77 77 77
BOBO 1 171\ -\ 17 77 99 77
$S/V_{domain}$ 0.2 180.2\ -\ 18 33 33 33
$R_{soma}$ 0.5 160.5\ -\ 16 55 33 66
$R_{MRsoma}$ 3 113\ -\ 11 77 44 88
$\eta_{s}oma$ 0.5 1000.5\ -\ 100 $17$ $25$ $13$
$S/V_{soma}$ 0.05 300.05\ -\ 30 $0.7$ 11 $0.6$
$S/V_{branch}$ 4 194\ -\ 19 77 88 66
$L_{branch}$ 3 1903\ -\ 190 $54$ $13$ $68$
$R_{branch}$ $\leq 0.04\ -\ 1.6$ $0.6$ $0.5$ $0.6$
$R_{MRbranch}$ 0.20 1.050.20\ -\ 1.05 $0.7$ $0.6$ $0.7$
$CV_{branch}$ 0.05 4.60.05\ -\ 4.6 $0.6$ $0.8$ $0.6$
$\mu OD_{branch}$ 0.05 0.600.05\ -\ 0.60 $0.25$ $0.23$ $0.26$
$R_{c}$ $\geq 1\ -\ 640$ $29$ $16$ $32$
$\tau_{branch}$ 0.43 0.950.43\ -\ 0.95 $0.77$ $0.75$ $0.78$
Table 3: Reference values for all the morphological features of neuronal and glial cells. The ranges and mean values obtained from the whole dataset investigated are reported, together with mean values for only neurons and glia. The $\geq$’ and $\leq$’ are used when the estimated value of the corresponding feature may be slightly (on average $20\%$) under- or over-estimated, respectively, given the known limitations of the approach used.
Cell Type Species $\kappa$ OD $(\tau_{1},\tau_{2},\tau_{3})$ FA adjFA
Mouse/Rat $1.4\pm 2.2$ $0.51\pm 0.22$ (0.13,0.31,0.56) $0.55\pm 0.22$ $0.33\pm 0.23$
Microglia Monkey $1.1\pm 0.6$ $0.50\pm 0.19$ (0.01,0.35,0.64) $0.76\pm 0.05$ $0.36\pm 0.17$
Human n.a. n.a. n.a. n.a. n.a.
Mouse/Rat $1.6\pm 1.3$ $0.43\pm 0.22$ (0.01,0.29,0.61) $0.63\pm 0.20$ $0.41\pm 0.22$
Astrocyte Monkey n.a. n.a. n.a. n.a. n.a.
Human n.a. n.a. n.a. n.a. n.a.
Mouse/Rat $6.0\pm 4.6$ $0.19\pm 0.19$ (0.03,0.15,0.82) $0.87\pm 0.15$ $0.76\pm 0.23$
Oligodendrocyte Monkey n.a. n.a. n.a. n.a. n.a.
Human n.a. n.a. n.a. n.a. n.a.
Mouse/Rat $1.6\pm 0.12$ $0.44\pm 0.20$ (0.09,0.25,0.66) $0.70\pm 0.16$ $0.52\pm 0.23$
Pyramidal Monkey $0.9\pm 0.6$ $0.57\pm 0.19$ (0.06,0.29,0.65) $0.72\pm 0.10$ $0.45\pm 0.23$
Human $1.1\pm 0.8$ $0.54\pm 0.20$ (0.14,0.27,0.59) $0.60\pm 0.14$ $0.43\pm 0.21$
Mouse/Rat $4.4\pm 2.8$ $0.18\pm 0.11$ (0.04,0.15,0.81) $0.87\pm 0.09$ $0.77\pm 0.16$
Granule Monkey n.a. n.a. n.a. n.a. n.a.
Human $3.7\pm 1.5$ $0.19\pm 0.08$ (0.05,0.16,0.78) $0.85\pm 0.07$ $0.75\pm 0.13$
Mouse/Rat $1.5\pm 0.7$ $0.42\pm 0.17$ (0.01,0.28,0.71) $0.80\pm 0.08$ $0.52\pm 0.19$
Purkinje Monkey n.a. n.a. n.a. n.a. n.a.
Human n.a. n.a. n.a. n.a. n.a.
Mouse/Rat $1.3\pm 0.8$ $0.49\pm 0.19$ (0.11,0.29,0.60) $0.62\pm 0.16$ $0.40\pm 0.20$
Basket Monkey n.a. n.a. n.a. n.a. n.a.
Human $1.2\pm 0.7$ $0.49\pm 0.20$ (0.12,0.34,0.54) $0.56\pm 0.13$ $0.28\pm 0.14$
Mouse/Rat $1.3\pm 0.9$ $0.49\pm 0.20$ (0.07,0.30,0.63) $0.69\pm 0.13$ $0.43\pm 0.21$
Gabaergic Monkey $3.2\pm 1.6$ $0.24\pm 0.15$ (0.07,0,20,0.73) $0.78\pm 0.13$ $0.64\pm 0.24$
Human $2.5\pm 1.4$ $0.29\pm 0.14$ (0.3,0.25,0.62) $0.64\pm 0.15$ $0.50\pm 0.15$
Mouse/Rat $1.1\pm 0.9$ $0.55\pm 0.21$ (0.08,0.33,0.59) $0.65\pm 0.12$ $0.33\pm 0.21$
Glutamatergic Monkey n.a. n.a. n.a. n.a. n.a.
Human n.a. n.a. n.a. n.a. n.a.
@@ 3.5 Distribution of some structural features of interest for dMRI modelling
The full distributions of values for some structural features of interest, obtained by merging together all the estimated values from all the species for each cell type, are shown in Fig.6. Given the increasing interest of the diffusion-based microstructural imaging community in estimating glia microstructure and the striking difference between the neuronal components of cerebral and cerebellar cortices (e.g., Purkinje cells only in cerebellum), we decided to group them into three classes: glia, cortical neurons and cerebellar neurons, and use three different colors to simplify the visualization of the results. The distributions for all the morphological features are reported in Supplementary Fig.3.
Fig.7 shows the distributions of intra-cellular and intra-branch residence times (grouped by general cell types) for two representative membrane permeabilities (low and high), as well as mean residence and exchange times as a function of different values of cell membrane permeability.
Figure 6: Distributions of soma radius, branch length, branch surface-to-volume ratio, and cellular surface-to-volume ratio (grouped by general cell types). Dashed lines in soma radius distribution plot indicate the effective MR radius for each general cell type (Glia $R_{MRsoma}=3.7\mu m$, Cortex neurons $R_{MRsoma}=10.4\mu m$, and Cerebellum neuron $R_{MRsoma}=7.1\mu m$
Figure 7: (A) Distributions of intra-cellular and intra-branch residence times (grouped by general cell types) for two membrane permeabilities (low and high). Residence time distributions were derived from surface-to-volume ratio distributions using permeability values of 2 and 20 $\mu m/s$ to represent low and high permeability, respectively. (B) Mean residence and exchange times as a function of membrane permeability. The mean values were calculated from the distributions of exchange times obtained for different surface-to-volume ratios and membrane permeabilities. See Section 2.2 for details.
@@ 3.6 Topological distance between cell types
The persistence maps of all the cell types for all species together are shown in Fig.8. These persistence maps are used for each species and each cell type to estimate cell-type specific topological distances, shown in Fig.9. Under bootstrapping of n=1000 iterations, it was found all distances were statistically significant between cell types, except basket and glutamatergic rodent cells with a p = 0.16.
In rodent cells, there is a notably high global topological distance between glial cells and neurons, indicating a fundamentally different topological organisation between them. This difference in topology becomes even more evident when considering the comparison along the same length scale. Since glial cells are significantly smaller in size, they inherently display a much smaller topological persistence. This size-related constraint underscores their distinct spatial and structural properties relative to neurons.
Within rodent neurons, granule cells stand out by exhibiting a higher topological distance compared to other neuronal types. This disparity is likely related to their characteristically low branch number, which reflects their simpler dendritic structures and reduced connectivity compared to more complex neurons.
In human cells, however, no significant or consistent trends in topological distance are observed, suggesting greater variability or less pronounced differences in cellular topology among the various neuronal cell types.
When examining hominid cells, which include both monkey and human samples, microglia demonstrate a higher topological distance compared to neurons. This observation reinforces the trend seen in rodent cells, further supporting the notion that glial cells maintain a distinct and conserved topological profile across different species.
In the cross-species comparison between rodent and hominid cells, microglia consistently show a high topological distance relative to neurons, mirroring the trend observed within each species. Additionally, the comparatively low topological distance for microglia across species suggest that their topological properties are consistent, pointing to a shared structural and functional organization in microglia across species.
@@ 4.1 A Quantitative View of Gray Matter Microstructure
[…]
One limitation of our analysis is that we did not include spines, boutons and glial leaflets, that can occupy up to $20\%$ of the GM volume [67]. Dendritic spines influence intracellular diffusion by increasing the surface-to-volume ratio and introducing structural complexity and compartmentalisation within dendrites [68, 69, 70]. Using simple modeling (see Supplementary Fig.5), we estimate that a spine density of 1 spine/µm increases the branch surface-to-volume ratio by $\approx 20–30\%$ in rodent neurons, leading to a $\approx 15–20\%$ reduction in residence time. At a density of 2 spines/µm, the surface-to-volume ratio increases by $\approx 50\%$, with a corresponding $\approx 35\%$ reduction in residence time.
Beyond these surface effects, spines act as distinct compartments: their narrow necks serve as diffusion barriers, restricting molecular exchange between the spine head and dendritic shaft and thereby altering diffusive dynamics ([69, 70]). Consequently, spine morphology represent an important morphological feature to be considered in models of molecular diffusion to accurately reflect in vivo conditions, but its in depth investigation is beyond the scope of this work.
In general, exchange cannot be completely ignored, nor does exchange fully average out the restriction effects of cell membranes, supporting including both neurite exchange and a soma compartment in a parsimonious biophysical model of diffusion in GM, as proposed by Olesen et al. [34] with the SANDIX model. There are however some caveats to bear in mind when interpreting the model estimates: (a) the soma apparent MR radius is unavoidably an overestimation of the true soma radius (Tab.2); (b) the compartmental signal fractions reflect, apart from volume, also unknown T2 and T1 weighting, which might differ among the compartments; (c) fast exchange across the neurite membrane’s could be problematic for models based on the Kärger’s model because it could violate the model’s assumptions (i.e. barrier-limited diffusion: $t_{ex}\gg d^{2}/D$, where $d$ is the neurite diameter and $D$ the molecular diffusivity; and $\delta\ll t_{ex}$, where $\delta$ is the gradient pulse duration of a single diffusion encoding dMRI measurement) and lead to biased estimates; (d) molecular diffusion in structural features such as dendritic spines and glial leaflets can lead to diffusion-mediated exchange mechanisms that occur on the same time scales of permeative exchange, making the interpretation of exchange measurements through MRI solely due to permeative exchange fundamentally wrong [68, 69, 70]; (e) using single diffusion encoding acquisitions it is impossible to disentangle restriction and exchange; more refined acquisitions could allow the separation and more accurate quantification of these two effects [64, 71, 69, 70]. The estimates of the model parameters should therefore only be taken as an indication of the true tissue features and validation against realistic numerical simulations (e.g., using the high resolution 3D exemplar meshes we provide for each cell type) and/or alternative measurements in controlled phantoms (e.g. biomimetic tissues and brain organoids) and/or post-mortem samples (e.g., optical, confocal or electron microscopy) remains essential.
@@ 4.2 General Implications for Biophysical Modelling
Here we provide a few illustrative examples demonstrating how our results can inform biophysical modelling. We discuss both general considerations and a representative in vivo acquisition case using parameters typical for clinical scanners.
@@
Assuming an intra-cellular diffusivity D=2 $\mu m^{2}/ms$ representative of water and 0.40 $\mu m^{2}/ms$ representative of intracellular brain metabolites, and considering a typical single diffusion encoding acquisition with gradient pulse duration $\delta<30\ ms$, gradient pulse separation $\Delta<70\ ms$ and diffusion time td<60mstd<60\ ms, we can infer from from our results:
@@
  • Impact of soma restriction. The impact of soma restriction is measurable when (from [72]) $5Dt_{d}\geq R_{soma}^{2}$ , that is for $t_{d}\geq 2.5\ ms$ for water and $t_{d}\geq 12.5\ ms$ for metabolites, given a soma radius $R_{soma}\approx 5\ \mu m$. These reference values become slightly longer if we consider the effective MR radius $R_{MRsoma}\approx 7\ \mu m$: $t_{d}\geq 5\ ms$ for water and $t_{d}\geq 24.5\ ms$ for metaboilites. Given the exemplar case of the in vivo acquisition, this suggests that soma restriction can have a measurable impact for both water and metabolites. This is supported by previous findings which demonstrate the impact of soma contribution on the measured dMRI signal in GM [32, 46]
  • Impact of diffusion-mediated exchange between soma and projections. Considering the total area of the soma surface covered by the openings towards cellular projections $A_{proj}\approx N_{proj}\pi R_{branch}^{2}$ and the soma volume $V_{soma}\approx\frac{4}{3}\pi R_{soma}^{3}$, the expected residence time within the soma is in first approximation $\tau_{i}^{soma}\approx\frac{V_{soma}}{4\sqrt{\frac{A_{proj}}{\pi}}D}=\frac{\pi R_{soma}^{3}}{3R_{branch}\sqrt{N_{proj}}D}$. Assuming that the soma to projection volume ratio is approximately 1:4 (from our data we estimated 1:4.1 for glia, 1:4.7 for cortical neurons and 1:4.4 for cerebellar neurons), the impact of diffusion-mediated exchange between soma and projection is measurable if $t_{d}\geq(\frac{1}{\tau_{i}^{soma}}+\frac{1}{\tau_{i}^{branch}})^{-1}=(\frac{1}{\tau_{i}^{soma}}+\frac{1}{3}\frac{1}{\tau_{i}^{soma}})^{-1}$, hence for $t_{d}\geq 31\ ms$ for water and $t_{d}\geq 155\ ms$ for metabolites. These estimates become longer if we consider the effective MR radii $R_{MRsoma}$ and $R_{MRbranch}$: $t_{d}\geq 73\ ms$ for water and $t_{d}\geq 364\ ms$ for metabolites. Given the exemplar case of the in vivo acquisition, this suggests that the impact of diffusion-mediated exchange between soma and projections is likely negligible or minimal for both water and metabolites.
@@
  • Impact of cellular domain restriction: given the size of the cellular domain, it will only significantly restrict molecular diffusion when $5Dt_{d}\geq R_{domain}^{2}$, that is for $t_{d}\geq 360\ ms$ for water and $t_{d}\geq 1800\ ms$ for metabolites, given $R_{domain}\geq 60\ \mu m$, which is far longer than typically used diffusion times $t_{d}$. Given the exemplar case of the in vivo acquisition, this suggests that the impact of cellular domain restriction is generally negligible for both water and metabolites.
@@
  • Impact of projections curvedness. The impact of curvedness can only be significant when (from [73], $2D\Delta,2D\delta\geq(<R_{c}>_{s})^{2}$, that is for $\Delta,\delta\geq 225\ ms$ for water and $\Delta,\delta\geq 1125\ ms$ for metabolites, given $R_{c}>_{s}\approx 30\ \mu m$, which is far longer than the values typically used. Given the exemplar case of the in vivo acquisition, this suggests that the impact of projections curvedness is generally negligible for both water and metabolites. This is further supported by [46].
@@
  • Impact of projection undulation: the values of $\mu OD_{branch}>_{s}$ estimated from the real cellular data match those simulated in [48], from which it can be concluded that undulation can have a measurable impact for both water and metabolites and can bias the estimation of projection radius in GM. However, the average branch radius $R_{branch}>_{s}\approx 0.6\ \mu m$ is far below the resolution limit of conventional water dMRI techniques [74].
@@
  • Impact of branching. Given the branch length $L_{branch}\approx 54\ \mu m$, the exchange between branches is negligible for $t_{d}<<L_{branch}^{2}/(2D)\approx 750\ ms$ for water and $t_{d}<<3750\ ms$ for metabolites, significantly longer than typical $t_{d}$ used. Given the exemplar case of the in vivo acquisition, this suggests that the impact of branching is generally negligible for both water and metabolites.Further supported by [44, 46].
@@
  • Impact of water permeative exchange between cellular projections and extra-cellular space. The average intra-branch residence times given the estimated $S/V_{branch}$ range from $\approx 10\ ms$ to $\approx 98\ ms$ for neurons and from $\approx 9\ ms$ to $\approx 95\ ms$ for glial cells (these are purely based on projections morphology and exemplar membrane permeability 2-20 $\mu m/s$; we do not account for active transport nor water channels). The corresponding exchange times (considering 30% in volume occupied by extra-cellular space) range from $\approx 3\ ms$ to $\approx 30\ ms$ for neurons and from $\approx 3\ ms$ to $\approx 29\ ms$ for glial cells. The permeative exchange is thus negligible only for $t_{d}<30\ ms$ for both neuronal and glial projections; which is not satisfied in conventional water dMRI applications. Given the exemplar case of the in vivo acquisition, this suggests that the impact of water permeative exchange between cellular projections and extra-cellular space is measurable. Further supported by [33, 46, 66].
@@
  • Impact of projections orientation dispersion. Cells with highly oriented and polarized projections, such as Purkinje and granule cells have high FA (>0.50>0.50) and low orientation dispersion (<0.25<0.25); while the projections of most of the glial cells and other neuronal cells have FA (<0.50<0.50) and high orientation dispersion (>0.25>0.25). Given the exemplar case of the in vivo acquisition, this suggests that the impact of projections orientation dispersion is measurable. This supports growing evidence that FA, e.g., from DTI, and orientation dispersion estimates, e.g. from NODDI, can discriminate different cytoarchitectural domains (or layers) in hippocampus, cortex and cerebellum [75, 76, 77, 78, 79].
  • Glia topology is significantly different from neurons topology in rodents. The overall cellular topology can be conceptualized as a network of interconnected compartments, comprising the soma and its cellular projections. This large-scale organization is as critical for intracellular diffusion as the finer structural details. In neural cells, projections form a branched tree-like architecture, where each branching point functions as a diffusion junction. As a result, molecular diffusion - and consequently the dMRI signa - is strongly constrained by this branching topology. Biophysical models of the intracellular dMRI signal must therefore explicitly account for these topological features. Results in Fig.9 show that glial cells in rodents have a very different topology from neuronal cells, suggesting that it may be possible (with appropriate modelling) to disentangle glial and neuronal contributions to the measured dMRI signal. This is further supported by[80].